jhonyy9 (jhonyy9):

- true or false ? - in plane of naturale numbers 1/2 is a part of every primes greater than 2 .

jhonyy9 (jhonyy9):

p=2n+1 by definition - divide both sides by 2 p/2 = 2n/2 +1/2 = n + 1/2 3/2 =(2*1+1)/2 =3/2 -- for n=1 5/2 =(2*2+1)/2 =5/2 -- for n=2 . . . .

OpenStudy (anonymous):

hi @jhonyy9 hi @KatrinaKaif

OpenStudy (anonymous):

books.google.co.in/books?isbn=0691118809... go to this !

jhonyy9 (jhonyy9):

what will be there ?

jhonyy9 (jhonyy9):

please !

OpenStudy (anonymous):

true..

jhonyy9 (jhonyy9):

- ok but can being this statement proven ? ...how ?

OpenStudy (anonymous):

logic!!

jhonyy9 (jhonyy9):

- ..., yes is true but i think that you know this too that in matamatics aiways every statement need to be proven ...for ... - so ... ???

OpenStudy (mani_jha):

What do you mean by '1/2 is a part of every primes'? What does it mean for a number to be a part of 'a prime'?

jhonyy9 (jhonyy9):

- so the same like in case of Ryemann's hypothesys in plane of complex numbers where this say that always for every primes there the real part is 1/2

jhonyy9 (jhonyy9):

now this is clearly ?

OpenStudy (mani_jha):

Ya, But p=2n+1 is not a prime number when n=4,7,10etcSo it is probably false.

jhonyy9 (jhonyy9):

read you please my first answer !

jhonyy9 (jhonyy9):

by definition every primes can be writeing in the form of 2n+1

OpenStudy (anonymous):

I'm just curious, what area of maths is this? I am doing elementary number theory at the moment, and I don't know what you mean by "1/2 is part of every prime number greater than 2". In what sense is it part of it? By the plane of natural numbers, do you mean that you have two dimensions, where each dimension takes input from natural numbers, so that points on that plane are (a,b) where a and b are both natural numbers, so that maybe (a,b) has meaning a/b or something? Sorry for my unhelpful reply, perhaps! You are probably doing an area of maths I don't know about...

OpenStudy (anonymous):

True

jhonyy9 (jhonyy9):

,,MusicalAnimal" - what is true ? and how ? please ...- if possibile ,can you prove it too ?

OpenStudy (anonymous):

false

OpenStudy (anonymous):

3 is prime no and its factorization does not include 2

jhonyy9 (jhonyy9):

- this statement not was 2 ,look it there is 1/2

jhonyy9 (jhonyy9):

- so like 1 + 1/2 = 3/2 where 3 is prime